Jones Manufacturing purchased $10,000 of merchandise inventory on account from a vendor and paid a $500 freight bill. The credit terms are 2/10 or n/30. Because some of the merchandise was not needed, Jones Manufacturing returned $2,000 the same day. Jones Manufacturing uses the perpetual inventory system and made payment for the merchandise, less the return, within the discount period. What is the final cost of the merchandise inventory for Jones Manufacturing from this purchase?

A: 8,000 B: $7,840 C: 8,340 D: $10,094
Answer #1

SOLUTION

The final cost of the merchandise inventory for Jones Manufacturing from this purchase is $8,340 thus Option C is correct.

Net purchases = Purchases - Purchase returns

= $10,000 - $2,000

= $8,000

Net purchases payment after discount = 8,000 - (8,000*2%)

= 8,000 - 160 = 7,840

Final payment = Net purchases payment after discount + Freight charges

= $7,840 + 500 = $8,340